OpenStudy (anonymous):

What are three irrational numbers between 4 and5

geerky42 (geerky42):

4 = √4^2 = √16 5 = √5^2 = √25 so anything between √16 to √25, such as √17, √18, √19, etc until √25

geerky42 (geerky42):

Is that clear? Do you understand my approach?